OpenStudy (anonymous):

Write a formula for the general term (the nth term) of the geometric sequence. 1/2 ; -(1/10) ; 1/50 ; -(1/250), ..... A. an = (1/2)n - 1(3/5) B. an = 1/2 - 1/5(n - 1) C. an = 1/2(-1/5)n - 1 D. an = 1/5(-1/2)n - 1

OpenStudy (anonymous):

Alright, I sort of understand.

OpenStudy (anonymous):

1/25, right?

OpenStudy (anonymous):

Good.. Can you do it for second one now??

OpenStudy (anonymous):

It's also 1/25, is it not?

OpenStudy (anonymous):

Oh no!

OpenStudy (anonymous):

And to think, it was making sense. -laughs-

OpenStudy (anonymous):

Ha ha ha...

OpenStudy (anonymous):

See, simply find r by dividing second term with first term that is 1/2 with -1/10 ...

OpenStudy (anonymous):

Divide -1/10 by 1/2..

OpenStudy (anonymous):

Which is -1/5, so... Looking at my options.

OpenStudy (anonymous):

Yes it is -1/5..

OpenStudy (anonymous):

We must have to find another though, right?

OpenStudy (anonymous):

\[\large a_n = a \cdot r^{n-1} \implies \frac{1}{2} \cdot (-\frac{1}{5})^{n-1}\]

OpenStudy (anonymous):

What choice is this??

OpenStudy (anonymous):

*which

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

Good...

OpenStudy (anonymous):

That was depressingly easy. Thank you!

OpenStudy (anonymous):

For showing exponent or power, please use "^" this sign..

OpenStudy (anonymous):

You are welcome dear...